### v2

# a case of Macau algorithms

Abstract
A proof-of-randomness (PoR) protocol is a fair and low energy-cost consensus mechanism for blockchains. Each network node of a blockchain may use a true random number generator (TRNG) and hash algorism to fulfil the PoR protocol. In this whitepaper, we give the consensus mechanism of the PoR protocol, and classify it into a new kind of randomized algorithms called Macau. The PoR protocol could generate a blockchain without any competition of computing power or stake of cryptocurrency. Besides, we give some advantages of integrating quantum random number generator (QRNG) chips into hardware wallets, and also discuss the way to cooperate with quantum key distribution (QKD) technology.
I introduction
Blockchian is a decentralized data recording technology of a network. All network nodes are distributed in space, while blocks are distributed in time-order with a chain connection. The current block usually records its generation timestamp, the nonce of the node who generate the current block, the hash value of the last block’s data, and all transactions in the network after last block. Consensus algorithms play the central role of generating blockchains. The first blockchain, bitcoinbitcoin , use proof-of-work (PoW) consensus to select a network node for a new block. The PoW requires a hashing computer to calculate a hash value with very small probability. For example, a 256-bit hash value with all zeros in highest 7070~{}7070 bits. Such method consumes more and more computing power, which means a waste of massive electricity energy. Recently, Ethereum ethereum replace its PoW by proof-of-stake (PoS) method to save more than $99\%$ power, however, the fairness of the PoS is questioned.
In this whitepaper, we present a physically fair and low-energy cost consensus algorithm called proof-of-randomness (PoR) protocol. The PoR requires true random number generator (TRNG) hardware on each network node. The fairness comes form the physically true random number, and the main power consumption is from a TRNG chip which is usually less than 10mW on each node. Since PoR depends on multiple random number inputs to get final output, it is not include in either Monte Carlo or Las Vegas randomized algorithms. We define such new kind of randomized algorithms as Macau, comparatively, and discuss their general features.
II generation and testing of true random numbers
Before running PoR protocol, every network node should acquire an array of binary true random numbers form a TRNG hardware. For stationary computers, such as servers or desktop PCs, a big TRNG device or a PCI card with TRNG chips could work. However, for mobile computers such as laptops, smart phones or tablet PC, a mobile USB device with quantum random number generator (QRNG) chips is a good choice. Since quantum randomness is the fundamental true randomness of the nature (Einstein called it ’God plays dice’), QRNGs perform much better than classical TRNGs in complex environment and also has much lower energy consumption and much smaller size than classical TRNG chips IDQ-whitepaper . A mobile hardware wallet with QRNG chip is a suitable solution for blockchain network nodes.
We can test the quality of random numbers by embedding testsuits into blockchain softwares, e.g., NIST SP 800-22 randomness testsuitNIST-test . This step is necessary to forbid any pseudo random number generator (PRNG) since any pseudo random number is indeed determined and not fair enough for PoR. Typically, it requires 256256~{}256256 Mb binary random numbers to finish the test. When a node’s QRNG (or TRNG) chip has output 256 Mb random number data and passed the test, the blockchain software could split the 256 Mb data into 256 pieces, for example, with each piece’s length at 1Mb. Then the blockchain software could get a random byte from the QRNG (or TRNG) chip again as a serial number to select out 1Mb random number date. We name it as $R_{n}(m)$ with the blockchain time-sequence number $m$ and the network node number $n$.
III proof-of-randomness protocol
When a network node finishes its true random number test, the selected 1 Mb random number $R(m,n)$ is calculated by SHA-256 to get a 256-bit hash value $H_{n}(m)$, which is then sent to other nodes together with this node’s digital signature.
In the case that the network is relatively small, assuming $N$ nodes, the synchronizing of data among all $N$ online nodes could be within several minutes. Then each node can receive every other online node’s 1Mb random number and every node would acquire the same sum of 1 Mb random numbers, which is
$S_{H}(m)=\sum_{n=1}^{N}H(m,n).$ (1)
This sum is independent to the order of receiving other nodes’ $R(m,n)$ on a node due to commutative law. Then every node calculate the hash value of $S_{H}(n)$ by SHA-256 and broadcast the result $HS(n)$ to all other nodes together with its digital signature. Here we call $HS(m)$ the second hash value in order to distinguish the first hash value $H_{n}(m)$ of every node. Since $HS(m)$ depends on all 1 Mb random numbers form every node, it is unpredictable. While the absolute difference value $D(m,n)=|HS(m)-H_{n}(m)|$ is also unpredictable, which prevents the attacks of using controllable random numbers form any node. A consensus of choosing the node with minimal or maximal $D(m,n)$ as the $mth$ block’s owner is the final step of PoR.
Figure 1: With proof-of-randomness (PoR), a block records owner’s 1Mb true random number instead of a ’nonce’ in PoW.
In the case that the network is relatively big, synchronizing of data among all nodes is not possible. It would probably happen in large public blockchains. Here we should not calculate the second hash value $HS(m)$ anymore since different nodes may get different values. Instead, every node should calculate the hash value of $H_{n}(m)+H(m-1)$ by SHA-256 and broadcast the result $H^{\prime}_{n}(m)$, where $H(m-1)$ is the prevhash value of the last block. Then every node calculate $D(m,n)=|H^{\prime}_{n}(m)-H_{n}(m)|$ and broadcast $D(m,n)$ together with its digital signature. A consensus of choosing the node with minimal or maximal $D(m,n)$ as the $mth$ block’s owner still works.
However, a kind of attack could be possible in such relatively big network, which is to try multiple generation of random numbers and select a much closer $R(m,n)$ to the target value (minimal or maximal $D(m,n)$). A node with more QRNG chips or with expensive TRNG devices could do the attack. In order to prevent such kind of drawback to PoW (with a competition of random number generating power instead of computing power), a node $n$ should broadcast its time-stamp of finishing the true random number test, $t_{1}(m,n)$, and its time-stamp of finishing $D(m,n)$, which is $t_{2}(m,n)$. A consensus of choosing the node with minimal $D(m,n)\Delta t(m,n)$ as the $mth$ block’s owner could be reasonable, where $\Delta t(m,n)=t_{2}(m,n)-t_{1}(m,n)$. If a node try to generate random numbers by multiple times, $\Delta t(m,n)$S would increase and reduce its possibility of get minimal $D(m,n)\Delta t(m,n)$. Besides, $\Delta t(m,n)$ should set a minimal value in order not to be attacked by forging timestamps.
IV Randomized algorithm: Macau
There are two kinds of widely used randomized algorithms. One is Monte Carlo, which use limited steps of random inputs to obtain a output with a high probability (usually close to one) to be the predetermined target value. PoS is an example of Monte Carlo algorithms. The other is Las Vegas, which use unlimited steps of random inputs to obtain a output that must be the predetermined target value. PoW is a very typical example of Las Vegas algorithms, where a hash value with  70 zeros in high bits is the predetermined target, and the number of random attempts is not limited until such hash value is calculated out in a network node.
PoR is a randomized algorithm with true random number inputs, while the output depends on the hash value of all inputs. Since the output is not predetermined, PoR can not be classified into either Monte Carlo or Las Vegas algorithms. Here we define a new kind of randomized algorithms called Macau to classify PoR. Comparing to Monte Carlo and Las Vegas, Macau algorithms use limited steps of random inputs to obtain a output that is not predetermined before any inputs.
Figure 2: Brief differences between Macau algorithm and Monte Carlo algorithm, as well as Las Vegas algorithm.
We give a general description of Macau algorithms by Eq.2
$T=F[\sum_{n=1}^{N}R(n)],$ (2)
where $T$ is the target value, $R(n)$ is the $nth$ random number input, and $F[x]$ is a function of the variable which is the sum of all input random numbers. In PoR, $F[x]$ is just the SHA-256 function of the sum of input random number’s hash values. Since both addition and multiplication have commutative law, the variable of $F[x]$ could be the product of all input random numbers too, which is
$T=F[\prod_{n=1}^{N}R(n)].$ (3)
Macau algorithms are different to Monte Carlo and Las Vegas, since both of them have predetermined target values $T_{p}$. The equation of either Monte Carlo or Las Vegas is like $T(m)=F[R(m)]$, where $R(m)$ is the $mth$ attempt of random number input, and $T(m)$ is the result of $mth$ attempt by function $F$. Monte Carlo algorithms limit the attempt times, where a $T(m)$ equals to $T_{p}$ could probably be found. Las Vegas algorithms never limit attempt times in order to find a $T(m)$ which equals to $T_{p}$.
V other advantages of using QRNGs in blockchains
Blockchains adopt elliptic curve cryptograph (ECC) as the main encryption and decryption method. The ellipse curve equation is
$y^{2}=ax^{3}+bx+c,$ (4)
while most of current blockchains use a simplified elliptic curve equation, $y^{2}=x^{3}+7$.
There are three kind of random numbers in ECC. The first one is a base-point on the elliptic curve, $G$, which starts a half-line to cross the elliptic curve. The second one is the private key, $k$, which is the reflection times of the half-line on the elliptic curve. The public key $K$ is the final point of the half-line on the elliptic curve determined by $G$ and $k$, and expressed by a ”multiplication” defined on the curve, which is $K=G*k$. Since it is extremely difficult to calculate $k$ with public $G$ and $K$ values by a classical computer, the ECC is secure to all classical computers now.
The third random number is $R$, which encrypt the data $D$ together with the public key $K$. The encryption equations are
$E_{1}=K*R+D$ (5)
$E_{2}=G*R,$ (6)
while the decryption equation with private key $k$ is
$E_{1}-E_{2}*k=K*R+D-E_{2}=G*R.$ (7)
Here we can see when $R$ is not safe, e.g. a pseudo random number that could be attacked by stealing its short seed, an attacker could obtain $D$ with $D=E_{1}-K*R$ since both $K$ and $E_{1}$ are public.
When QRNGs are evolved in network nodes, both $k$ and $R$ can use quantum random numbers to prevent any crack by computers. It makes $E_{1}$ much safer to travel over the network. Therefore, a hardware wallet with a QRNG could be an upgrade of ECC’s security, which is also an upgrade of blockchain’s security.
Besides, QRNG chips in hardware wallets could provide quantum keys to encrypt their private keys in order to recover them. For current hardware wallets, when a wallet is lost, the private key is lost too. Then all the cryptocurrency bounding to this private key would be lost forever. When a hardware wallet use a quantum key $q$ from a QRNG chip to encrypt its private key $k$, the encrypted data $E$ could be sent out the wallet and stored in a disk or on cloud serve. While the quantum key could be stored in an off-line memory device which is apart from $E$. If a hardware wallet is lost, one can recover its private key by decrypting $E$ with $q$. This is another advantage of hardware wallets with QRNG chips.
VI in cooperation with quantum key distribution (QKD)
Since ECC is vulnerable to general quantum computers, blockchains should upgrade their cryptograph to keep the future security when a general quantum computer with millions of qubits comes to reality. Post-quantum cryptograph (PQC) is a natural solution which cost a general quantum computer much longer time to break than breaking RSA-2048 and ECC. However, it can not guarantee that a future quantum algorithm which breaks the PQC in a very short time never exist.
Quantum key distribution (QKD) is the ultimate solution of preventing the attack form quantum computers QKD . In QKD, all the keys are quantum random numbers, and any interception of key distribution would break the keys. Therefore, QKD with one-time encryption is unconditional secure to any algorithm. However, QKD networks rely on optical fibers (or even satellites like Mocius) which restrict QKD terminals in cloud servers. If every node of a network have a server together with a QKD terminal, one can use one-time keys form QKD network to do peer-to-peer authentication between any two nodes.
However, QKD terminals can not be used in mobile networks since there is no single-photon technology in radio-frequency or microwave band. Here we could use QRNG chips in a mobile node to do an off-line local distribution of quantum keys to any server node with QKD terminals. Then an on-line key update could work when current keys encrypting next keys (KEK) for communication. This method could work in all mobile networks since it is indeed a classical communication encrypted by quantum keys. With this method being applied on mobile networks, and QKD being applied on fiber networks, a full quantum-safe network is feasible.
VII conclusions
In this white paper, we present a proof-of-randomness (PoR) protocol. Such protocol could give a fair and low energy-cost consensus mechanism to blockchains with true random number generators (TRNG) and hash algorithm. Since low-cost quantum number generators (QRNG) make the true random number resource more accessible to common users, this PoR protocol could popularize the blockchain technology, especially public blockchains for Web3 applications. We classify PoR to a new kind of randomized algorithms called Macau algorithms in order to distinguish PoR form widely-used Monte Carlo and Las Vegas algorithms.
Besides, a hardware wallet with a QRNG chip could not only give out quantum random numbers to do the PoR protocol, but also make elliptic curve cryptograph ECC much safer by replacing all pseudo random numbers with quantum random numbers. Even more, a quantum random number could be a quantum key to encrypt the private key in the wallet, then a backup of the encrypted private key outside a hardware wallet is possible..
Finally, QRNG chips could cooperate with quantum key distribution (QKD) in order to prevent any attack form general quantum computers to blockchains ultimately in the future.
VIII acknowledgement
As a MiraclePlus alumni company (F22), we thank MP for its accelerator program and founders’ community.